# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 3.1 Configuration space and full action
We now summarise the full configuration space and the action of the dynamical–lattice regulator. The basic kinematical variables are:
@@
We denote the full configuration collectively by $(x,U,\Phi)$. The geometry degrees of freedom $x$ are governed by a local $\mathrm{E}(d)$–invariant action $S_{x}[x]$ as in Section 2.3, while the gauge and matter fields are coupled to $x$ through a local field action $S_{\mathrm{fields}}[x,U,\Phi]$. The total action is
@@
For complete precision one first defines the functional integral in finite volume, for instance on a time-open cylinder $\Lambda_{T,L}\subset\mathbb{Z}^{d}$ (with the first lattice direction singled out as index time), where all measures are ordinary finite products. The infinite–lattice notation used below is then understood as the limit $\Lambda_{T,L}\nearrow\mathbb{Z}^{d}$ on local observables. Global translation/rotation zero modes of $x$ may be fixed (or quotiented) in finite volume; this does not affect the local structure of the actions below nor the reflection positivity statements proved later.
@@
where $Dx$ is the flat measure in $\mathcal{X}_{\mathrm{adm}}$, $DU$ is the product Haar measure over link variables, and $D\Phi$ is the standard product measure over matter fields (Lebesgue for bosons and the usual Grassmann measure for fermions).
We now present a representative class of gauge and matter actions on a dynamical mesh. In all cases the fields live on the abstract lattice $\Lambda$, and the interactions have the usual nearest–neighbour stencil, with couplings dressed by local geometry–dependent weights. For later Osterwalder–Schrader (OS) reflection-positivity arguments, one subtlety is important: the local metric and volume constructed from a forward corner frame are intrinsically biased with respect to the time direction. Rather than modifying the finite-difference stencils, we remove this bias by action averaging: we define two local geometric dressings, exchanged by the OS reflection, and average the resulting actions. This produces a manifestly reflection invariant field action while preserving locality and gauge invariance.
@@
Prototype actions and $\theta$–averaging. Recall from (2.6)-(2.8) the forward and backward frames $E(n)$ and $E^{\theta}\!(n)$, as well as the local metric $g(n)$ and volume $V(n)$ computed from $E(n)$. We first define “prototype” field actions $S^{\circ}[x,U,\Phi]$ by coupling the usual nearest-neighbour stencils to $g$ and $V$. To remove the forward-time bias, we then define the $\theta$-partner action $S^{\theta}[x,U,\Phi]$ by the same formulas but with every occurrence of a direction index $\mu$ replaced by $\theta(\mu)$ and with the corner frame replaced by its $\theta$–version. Averaging the two yields the field action
$S_{\mathrm{fields}}[x,U,\Phi]:=\tfrac{1}{2}\bigl(S^{\circ}[x,U,\Phi]+S^{\theta}[x,U,\Phi]\bigr).$ (3.6)
@@
Scalar field. Let $\phi:\Lambda\to\mathbb{C}^{m}$ be a complex scalar multiplet transforming in some unitary representation of $G$. For $\mu\in\Delta$ define the gauge-covariant signed forward difference
@@
where $\langle a,b\rangle=a^{\dagger}b$ and $W_{\mathrm{loc}}$ is a local gauge-invariant potential such as $W_{\mathrm{loc}}(\phi)=m_{0}^{2}|\phi|^{2}+\lambda_{0}|\phi|^{4}$ (or, for multiplets, a function of $\langle\phi,\phi\rangle$). We define the prototype scalar action $S_{\phi}^{\circ}=S_{\phi,\mathrm{kin}}^{\circ}+S_{\phi,\mathrm{pot}}^{\circ}$ and then set
@@
To reiterate the definition of $S_{\phi}^{\theta}$, it is defined by replacing $(g,V)$ with $(g^{\theta},V^{\theta})$ and applying $\mu\mapsto\theta(\mu)$ to the direction labels appearing in (3.8), so that only the time-direction differences are reversed. On a regular embedding x(n)=anx(n)=a\,n, one has $E(n)=E^{\theta}\!(n)=aI$, hence $S_{\phi}$ reduces to the standard lattice scalar action.
@@
where the weight $w^{\mu\nu}(n;x)>0$ depends on the embedding only through local Euclidean invariants built from $(g(n),V(n))$ (equivalently from $E(n)$). Two natural choices are
@@
Fermions. Fermionic actions on a dynamical lattice can be constructed in close analogy with the standard hypercubic case, with the difference that the discrete Dirac operator must be built from the local geometry induced by the embedding $x$. Let $\psi(n)$ be a lattice Dirac spinor in a representation of $G$, and let $\bar{\psi}(n)$ denote the Dirac adjoint. We fix constant Euclidean gamma matrices $\{\gamma_{a}\}_{a=1}^{d}$ satisfying $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$, and use the inverse frame to contract abstract finite differences into target-space directions. Define
@@
We define $S_{f}^{\theta}$ as usual, which in this case amounts to replacing $V(n)$ by $V^{\theta}(n)$ and $\Gamma_{\mu}(n)$ by $\Gamma_{\mu}^{\theta}(n)$, and set
@@
On a regular embedding x(n)=anx(n)=a\,n, we have $\Gamma_{\mu}(n)=\Gamma_{\mu}^{\theta}(n)=\gamma_{\mu}/a$, and so (3.18) reduces to the usual naive lattice Dirac operator. As the standard lattice is among the possible embeddings $x(n)$, we do not expect the dynamical geometry alone to remove species doubling. Standard remedies (Wilson, staggered, overlap, domain-wall) can nevertheless be implemented by replacing the corresponding finite-difference operators and Wilson terms by their geometric analogues built from the same local data $(E,g,V)$ and then applying the $\theta$–averaging prescription (3.6).
@@ 3.2 Euclidean invariance, orientation, and local isotropy
We now record the global Euclidean symmetries of the action and make explicit where a restriction to proper motions enters. Recall that the Euclidean group is the semidirect product
$\mathrm{E}(d):=O(d)\ltimes\mathbb{R}^{d},$ (3.22)
with multiplication $(R,b)(R^{\prime},b^{\prime})=(RR^{\prime},\,b+Rb^{\prime})$. We write $\mathrm{SE}(d):=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$ for the subgroup of proper Euclidean motions.
@@
where $\rho$ is the appropriate finite-dimensional representation of $O(d)$ on the matter multiplet (trivial for scalars; spinorial for Dirac fields, discussed below). Since $b$ is a translation in physical space, it acts trivially on internal indices.
Then the edge vectors transform as
@@
To verify global Euclidean invariance for the field action, it suffices to only deal with the prototype part $S^{\circ}$. Indeed, under $x(n)\mapsto Rx(n)+b$, the forward and backward frames transform as $E(n)\mapsto RE(n)$ and $E^{\theta}\!(n)\mapsto RE^{\theta}\!(n)$, so any coefficient built from Euclidean invariants is unchanged; the relabeling $\mu\mapsto\theta(\mu)$ is purely combinatorial. Therefore $S^{\theta}$ inherits the same $\mathrm{SE}(d)$-invariance as $S^{\circ}$, and so does the averaged action $\tfrac{1}{2}(S^{\circ}+S^{\theta})$.
@@
The potential term $S_{\phi,\mathrm{pot}}^{\circ}=\sum_{n}V(n)W_{\mathrm{loc}}(\phi(n))$ is also invariant provided $W_{\mathrm{loc}}$ is an O(d)O(d)–scalar in field space (e.g. $m_{0}^{2}|\phi|^{2}+\lambda_{0}|\phi|^{4}$).
Gauge actions. For the weighted Wilson–type plaquette action (3.13) the gauge field $U$ is unchanged by $(R,b)\in\mathrm{E}(d)$, so the traces are invariant. Therefore $S_{g}$ is $E(d)$-invariant provided each weight $w^{\mu\nu}(n;x)$ is built from Euclidean invariants of the local geometry (equivalently, from $g_{\mu\nu}(n)$, $g^{\mu\nu}(n)$, $V(n)$, and Gram minors). In particular this holds for both weight choices introduced in (3.14) and (3.15). (As in the continuum, weights built from oriented pseudoscalars would be sensitive to improper rotations.)
Fermionic actions. For fermions, the clean symmetry statement depends on whether one wishes to implement reflections on spinors. We therefore take $\mathrm{SE}(d)=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$ as the baseline continuous symmetry group when Dirac fields are present; extending to improper rotations requires the Pin group. Independently, the admissible set $\mathcal{X}_{\mathrm{adm}}$ typically selects an orientation sector, so the $x$-measure is in any case only $\mathrm{SE}(d)$-invariant.
@@
Under $x^{\prime}(n)=Rx(n)+b$ the local frame matrix transforms by $E^{\prime}(n)=R\,E(n)$, hence the geometry-dependent matrices $\Gamma_{\mu}(n)=\sum_{a}\gamma_{a}(E(n)^{-1})_{a\mu}$ transform by conjugation:
$\begin{split}\widehat{R}\Gamma_{\mu}(n)\widehat{R}^{-1}&=\sum_{a}\widehat{R}\gamma_{a}\widehat{R}^{-1}(E(n)^{-1})^{a}{}_{\mu}=\sum_{a}R_{a}{}^{b}\gamma_{b}\,(E(n)^{-1})^{a}{}_{\mu}\\ &=\sum_{a}\gamma_{b}\,(E(n)^{-1}R^{-1})^{b}{}_{\mu}=\Gamma_{\mu}^{\prime}(n).\end{split}$ (3.35)
Since the gauge links $U_{\mu}(n)$ are invariant, the “naive” Dirac operator transforms covariantly,
$(D_{\mathrm{naive}}^{\circ}\psi^{\prime})(n)=\widehat{R}(D_{\mathrm{naive}}^{\circ}\psi)(n),$ (3.36)
and using $V^{\prime}(n)=V(n)$ we conclude invariance of the fermion action:
@@
Local isotropy via twisting. It is useful to separate two distinct notions that could easily be conflated under the phrase “rotationally invariant regulator.” The first is an exact global covariance statement as we discussed so far in this subsection. However, this global covariance alone is not the “soul” of DLR. In the degenerate regime where $\mathcal{X}_{\mathrm{adm}}(a)$ nearly collapses to the rigid orbit of the regular embedding (so that $x(n)\approx aRn+b$), global covariance remains true, yet there is essentially no local twisting and no meaningful suppression of anisotropy beyond what an ordinary lattice already achieves as $a\to 0$. Thus global covariance is a necessary formal property but not a sufficient mechanism for improved isotropy at finite lattice spacing.
This leads us to the the second notion, a local isotropy property, which is the actual mechanism by which DLR can suppress lattice anisotropies. The relevant feature is that the embedding fluctuates in a way that mixes local orientations on short length scales. To quantify this we extract from the embedding a local orientation field. Define the dimensionless forward frame $\widetilde{E}(n):=a^{-1}E(n)$. By admissibility $\widetilde{E}(n)$ is invertible, hence it has a polar decomposition
@@
Note that we have $\det\widetilde{E}(n)>0$ by admissbility, so the polar factor $Q(n)$ lies in $\mathrm{SO}(d)$. We interpret $Q(n)$ as the local orientation and $S(n)$ as the local shape.
@@
These observables distinguish rigid orbit averaging from genuine local twisting: if $x(n)=aRn+b$ is rigid then $Q(n)\equiv R$ and hence $R_{\mu}(n)\equiv I$ and $\mathcal{R}_{\mu\nu}(n)\equiv I$; conversely, broad fluctuations of $R_{\mu}$ (together with short-range correlations) indicate nontrivial local twisting.
@@
where $\langle AB\rangle_{c}:=\langle AB\rangle-\langle A\rangle\langle B\rangle$ denotes the connected correlator. The quantity (3.41) measures how rapidly the local frame rotations decorrelate with separation. In a near-rigid regime (dominated by global zero modes) one has $R_{\mu}(n)\approx I$ and the fluctuations of $R_{\mu}$ are small and long-range, whereas in a genuine twisting regime the distribution of $R_{\mu}(n)$ is broad and the connected correlator decays on a short correlation length $\xi_{\mathrm{twist}}$ (typically a few lattice spacings). This short-range mixing of local orientations is precisely the ingredient that can reduce axis-versus-diagonal (direction-dependent) cutoff artefacts in matter correlators at finite $a$. Note that $R_{\mu}$ and (3.41) are invariant under global Euclidean motions $x\mapsto Rx+b$, hence are insensitive to the fixing of global zero modes.

### v2

@@ 3.1 Configuration space and full action
We summarise the full configuration space and the action of the dynamical lattice regulator. The basic kinematical variables are:
@@
We denote the full configuration collectively by $(x,U,\Phi)$. The geometry degrees of freedom $x$ are governed by a local action $S_{x}[x]$ built from Euclidean invariants of the embedding data as in §2.3, while the gauge and matter fields are coupled to $x$ through a local field action $S_{\mathrm{fields}}[x,U,\Phi]$. The total action is
@@
For complete precision one first defines the functional integral in finite volume, for instance on a time-open cylinder $\Lambda_{T,L}\subset\mathbb{Z}^{d}$, with the first lattice direction singled out as index time, where all measures are ordinary finite products. The infinite–lattice notation used below is then understood as the limit $\Lambda_{T,L}\nearrow\mathbb{Z}^{d}$ on local observables. Global translation/rotation zero modes of $x$ may be fixed (or quotiented) in finite volume; this does not affect the local structure of the actions below nor the reflection positivity statements proved later.
@@
where $Dx$ is the flat measure in $\mathcal{X}_{\mathrm{adm}}$, $DU$ is the product Haar measure over link variables, and $D\Phi$ is the standard product measure over matter fields, which is the Lebesgue measure for bosons and the usual Grassmann measure for fermions.
We present a representative class of gauge and matter actions on a dynamical mesh. In all cases the fields live on the abstract lattice $\Lambda$, and the interactions have the usual nearest neighbour stencil, with couplings dressed by local geometry dependent weights. For later Osterwalder–Schrader (OS) reflection-positivity arguments, one subtlety is important: the local metric and volume constructed from a forward corner frame are intrinsically biased with respect to the time direction. Rather than modifying the finite-difference stencils, we remove this bias by action averaging: we define two local geometric dressings, exchanged by the OS reflection, and average the resulting actions. This produces a manifestly reflection invariant field action while preserving locality and gauge invariance.
@@
Prototype actions and $\theta$–averaging. Recall from (2.6)-(2.8) the advanced and retarded frames $E(n)$ and $E^{\theta}\!(n)$, as well as the local metric $g(n)$ and volume $V(n)$ computed from $E(n)$. We first define “prototype” field actions $S^{\circ}[x,U,\Phi]$ by coupling the usual nearest-neighbour stencils to $g$ and $V$. To remove the forward-time bias, we then define the $\theta$-partner action $S^{\theta}[x,U,\Phi]$ by the same formulas but with every occurrence of a direction index $\mu$ replaced by $\theta(\mu)$ and with the corner frame replaced by its $\theta$–version. Averaging the two yields the field action
$S_{\mathrm{fields}}[x,U,\Phi]=\tfrac{1}{2}\bigl(S^{\circ}[x,U,\Phi]+S^{\theta}[x,U,\Phi]\bigr).$ (3.6)
@@
More generally, one may analogously symmetrize with respect to reflections in any coordinate direction (or build the action directly from such symmetrized geometric data), which reduces directional bias and makes the couplings more isotropic with respect to the lattice axes.
Scalar fields. Let $\phi:\Lambda\to\mathbb{C}^{m}$ be a complex scalar multiplet transforming in some unitary representation of $G$. For $\mu\in\Delta$ define the gauge-covariant signed forward difference
@@
where $\langle a,b\rangle=a^{\dagger}b$ and $W_{\mathrm{loc}}$ is a local gauge-invariant potential such as
$W_{\mathrm{loc}}(\phi)=\frac{m_{0}^{2}}{2}\phi^{\dagger}\phi+\frac{\lambda_{0}}{4!}(\phi^{\dagger}\phi)^{2}.$ (3.10)
We define the prototype scalar action $S_{\phi}^{\circ}=S_{\phi,\mathrm{kin}}^{\circ}+S_{\phi,\mathrm{pot}}^{\circ}$ and then set
@@
To reiterate the definition of $S_{\phi}^{\theta}$, it is defined by replacing $(g,V)$ with $(g^{\theta},V^{\theta})$ and applying $\mu\mapsto\theta(\mu)$ to the direction labels appearing in (3.8), so that only the time-direction differences are reversed. On the regular embedding $x(n)=an$ one has $E(n)=aI$ and $E^{\theta}(n)=aR_{\theta}$ with $R_{\theta}=\mathrm{diag}(-1,1,\dots,1)$. In particular $g(n)=E(n)^{\mathsf{T}}E(n)=a^{2}I$ and $g^{\theta}(n)=(E^{\theta}(n))^{\mathsf{T}}E^{\theta}(n)=a^{2}I$. Hence the $\theta$-averaged scalar action reduces to the standard hypercubic action.
@@
where the weight $w^{\mu\nu}(n;x)>0$ depends on the embedding only through local Euclidean invariants built from $(g(n),V(n))$, or equivalently from $E(n)$. Two natural choices are
@@
Fermions. Fermionic actions on a dynamical lattice can be constructed in close analogy with the standard hypercubic case, with the difference that the discrete Dirac operator must be built from the local geometry induced by the embedding $x$. Let $\psi(n)$ and $\bar{\psi}(n)$ be lattice Dirac spinors (Grassmann fields) in a representation of $G$. Fix Euclidean gamma matrices $\{\gamma_{a}\}_{a=1}^{d}$ with $\{\gamma_{a},\gamma_{b}\}=2\delta_{ab}$, and define
@@
We define the $\theta$-partner by the same stencil but with the advanced/retarded geometric data interchanged, i.e.
[…]
and
$S_{f}^{\theta}[x,\bar{\psi},\psi,U]=\sum_{n\in\Lambda}V^{\theta}(n)\,\bar{\psi}(n)\,(D_{\mathrm{naive}}^{\theta}\psi)(n).$ (3.22)
Finally set
@@
On a regular embedding x(n)=anx(n)=a\,n, the forward and backward corner frames coincide (with our convention for $E^{\theta}$), so $\Gamma_{\mu}(n)=\Gamma_{\mu}^{\theta}(n)=\gamma_{\mu}/a$, and (3.19) reduces to the usual naive lattice Dirac operator. As the standard lattice is among the possible embeddings $x(n)$, we do not expect the dynamical geometry alone to remove species doubling. Standard remedies (Wilson, staggered, overlap, domain-wall) can nevertheless be implemented by replacing the corresponding finite-difference operators and Wilson terms by their geometric analogues built from the same local data $(E,g,V)$ and then applying the $\theta$–averaging prescription (3.6).
@@ 3.2 Euclidean invariance, orientation, and local isotropy
We record the global Euclidean symmetries of the action and make explicit where a restriction to proper motions enters. Recall that the Euclidean group is the semidirect product
$\mathrm{E}(d)=O(d)\ltimes\mathbb{R}^{d},$ (3.25)
with multiplication $(R,b)(R^{\prime},b^{\prime})=(RR^{\prime},\,b+Rb^{\prime})$. We write $\mathrm{SE}(d)=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$ for the subgroup of proper Euclidean motions.
@@
where $\rho$ is the appropriate finite-dimensional representation of $O(d)$ on the matter field (trivial for scalars; spinorial for Dirac fields, discussed below). Since $b$ is a translation in physical space, it acts trivially on internal indices.
The edge vectors transform as
@@
To verify global Euclidean invariance for the field action, it suffices to only deal with the prototype part $S^{\circ}$. Indeed, under $x(n)\mapsto Rx(n)+b$, the advanced and retarded frames transform as $E(n)\mapsto RE(n)$ and $E^{\theta}\!(n)\mapsto RE^{\theta}\!(n)$, so any coefficient built from Euclidean invariants is unchanged; the relabeling $\mu\mapsto\theta(\mu)$ is purely combinatorial. Therefore $S^{\theta}$ inherits the same $\mathrm{SE}(d)$-invariance as $S^{\circ}$, and so does the averaged action $\tfrac{1}{2}(S^{\circ}+S^{\theta})$.
@@
The potential term $S_{\phi,\mathrm{pot}}^{\circ}=\sum_{n}V(n)W_{\mathrm{loc}}(\phi(n))$ is also invariant provided $W_{\mathrm{loc}}$ is an O(d)O(d)–scalar in field space (e.g. $\frac{1}{2}m_{0}^{2}|\phi|^{2}+\frac{1}{4!}\lambda_{0}|\phi|^{4}$).
Gauge actions. For the weighted Wilson–type plaquette action (3.14) the gauge field $U$ is unchanged by $(R,b)\in\mathrm{E}(d)$, so the traces are invariant. Therefore $S_{g}$ is $E(d)$-invariant provided each weight $w^{\mu\nu}(n;x)$ is built from Euclidean invariants of the local geometry (equivalently, from $g_{\mu\nu}(n)$, $g^{\mu\nu}(n)$, $V(n)$, and Gram minors). In particular this holds for both weight choices introduced in (3.15) and (3.16). Note that as in the continuum, weights built from oriented pseudoscalars would be sensitive to improper rotations.
Fermionic actions. In the presence of Dirac fields we work with the proper Euclidean group $\mathrm{SE}(d)=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$, since implementing reflections on spinors would require the Pin group. Independently, our admissible set $\mathcal{X}_{\mathrm{adm}}$ fixes an orientation component, so the xx–measure is restricted to $\mathrm{SE}(d)$ rather than $\mathrm{E}(d)$.
@@
Under $x^{\prime}(n)=Rx(n)+b$ the local frame matrices transform by $E^{\prime}(n)=R\,E(n)$ and likewise $E^{\theta\,\prime}(n)=R\,E^{\theta}(n),$ since both are built from the same local edge vectors of the embedding. Consequently the geometry–dependent matrices $\Gamma_{\mu}(n)$ and $\Gamma_{\mu}^{\theta}(n)$ transform by conjugation:
$\widehat{R}\Gamma_{\mu}(n)\widehat{R}^{-1}=\sum_{a}\widehat{R}\gamma_{a}\widehat{R}^{-1}(E^{-1})_{a\mu}=\sum_{a,b}R_{a}{}^{b}\,\gamma_{b}\,(E^{-1})_{a\mu}=\sum_{b}\gamma_{b}\,(E^{\prime-1})_{b\mu}=\Gamma_{\mu}^{\prime}(n),$ (3.38)
and similarly $\widehat{R}\Gamma_{\mu}^{\theta}(n)\widehat{R}^{-1}=\Gamma_{\mu}^{\theta\,\prime}(n)$. Since the gauge links $U_{\mu}(n)$ are unchanged by Euclidean motions and act only on gauge indices, the mixed prototype naive Dirac operator (3.19) transforms covariantly:
$(D_{\mathrm{naive}}^{\circ}\psi^{\prime})(n)=\widehat{R}(D_{\mathrm{naive}}^{\circ}\psi)(n).$ (3.39)
Using also $V^{\prime}(n)=V(n)$ (and $V^{\theta\,\prime}(n)=V^{\theta}(n)$) we conclude invariance of the prototype action:
@@
Local isotropy via twisting. It is useful to separate two notions that are sometimes conflated under the phrase “rotationally invariant regulator.” The first is the exact global covariance of the partition function under $\mathrm{SE}(d)$, which follows formally because the action depends on the embedding only through Euclidean invariants. However, this formal covariance is not, by itself, the operative mechanism at finite cutoff: it would remain true even if the geometry measure were concentrated near a rigid $\mathrm{SE}(d)$–orbit $x(n)\approx aRn+b$, in which case typical configurations exhibit essentially no local orientation mixing and one recovers the usual hypercubic anisotropies. What matters is therefore how global $\mathrm{SE}(d)$ covariance is realised in typical configurations after fixing/quotienting the global zero modes.
The second notion is a local isotropy (or local orientation–mixing) property. In the short-range geometry regime (SR) required for a local Symanzik description after integrating out geometry, the geometry sector cannot realise $\mathrm{SE}(d)$ covariance “rigidly” through long-range geometric modes; rather, global covariance must be implemented through short-range fluctuations of local frames. We make this precise by extracting from the embedding a local orientation field. Define the dimensionless forward frame $\widetilde{E}(n)=a^{-1}E(n)$. By admissibility $\widetilde{E}(n)$ is invertible, hence it has a polar decomposition
@@
Note that $\det\widetilde{E}(n)>0$ by admissibility, so the polar factor $Q(n)$ lies in $\mathrm{SO}(d)$. We interpret $Q(n)$ as the local orientation and $S(n)$ as the local shape.
@@
These observables distinguish rigid orbit averaging from genuine local twisting: if $x(n)=aRn+b$ is rigid then $Q(n)\equiv R$ and hence $R_{\mu}(n)\equiv I$ and $\mathcal{R}_{\mu\nu}(n)\equiv I$; conversely, broad fluctuations of $R_{\mu}$ together with short-range decorrelation indicate nontrivial local twisting.
@@
where $\langle AB\rangle_{c}=\langle AB\rangle-\langle A\rangle\langle B\rangle$ denotes the connected correlator. The quantity (3.44) measures how rapidly the local frame rotations decorrelate with separation. In a near-rigid regime (dominated by global zero modes) one has $R_{\mu}(n)\approx I$ and the fluctuations of $R_{\mu}$ are small and long-range, whereas in a genuine twisting regime the distribution of $R_{\mu}(n)$ is broad and the connected correlator decays on a short correlation length $\xi_{\mathrm{twist}}$ (typically a few lattice spacings). This short-range mixing of local orientations is precisely the ingredient that can reduce axis-versus-diagonal (direction-dependent) cutoff artefacts in matter correlators at finite $a$.